# Understanding Multimodal LLMs: the Mechanistic Interpretability of Llava in Visual Question Answering

### v1

@@ Abstract
Understanding the mechanisms behind Large Language Models (LLMs) is crucial for designing improved models and strategies. While recent studies have yielded valuable insights into the mechanisms of textual LLMs, the mechanisms of Multi-modal Large Language Models (MLLMs) remain underexplored. In this paper, we apply mechanistic interpretability methods to analyze the visual question answering (VQA) mechanisms in the first MLLM, Llava. We compare the mechanisms between VQA and textual QA (TQA) in color answering tasks and find that: a) VQA exhibits a mechanism similar to the in-context learning mechanism observed in TQA; b) the visual features exhibit significant interpretability when projecting the visual embeddings into the embedding space; and c) Llava enhances the existing capabilities of the corresponding textual LLM Vicuna during visual instruction tuning. Based on these findings, we develop an interpretability tool to help users and researchers identify important visual locations for final predictions, aiding in the understanding of visual hallucination. Our method demonstrates faster and more effective results compared to existing interpretability approaches. Code: https://github.com/zepingyu0512/llava-mechanism
@@ 3.2 Mechanism Exploration of Textual QA
Evidence c). Following Dar et al. (2022), we project the color position’s layer input vector $h_{p}^{l-1}$ into the unembedding space and calculate the MRR score for the animal tokens [A] and [A1]. In S0, the MRR for [A] is 0.756, while for [A1], it is 0.001, as shown in Figure 3(b). The logit difference between [A] and [A1] at the color position is 0.32. In S1, the MRR for [A] is 0.002, the MRR for [A1] is 0.715, and the logit difference between [A1] and [A] is 1.70. These scores demonstrate that the color position’s layer input vector in the most important heads encodes the animal features from its context.
@@ 3.3 Mechanism Exploration of Visual QA
Evidence c) and d). We project the layer inputs of the top 20 important positions into the unembedding space and compute the MRR scores and logit differences between the correct animal and a different animal. The correct animal’s MRR score is 0.318, while the other animal’s MRR score is 0.0004, as shown in Figure 4(b). The logit difference is 1.53, confirming that the important heads’ layer inputs contain crucial information about the animals. Furthermore, when the animal in the question is replaced with another animal, the attention score at the top 20 positions drops significantly from 0.807 to 0.564 (see Figure 4c), indicating that the last position encodes information about the question.
@@
Evidence e). Difference between VQA and TQA: Information Contained in Visual Embeddings. A key difference between Llava’s VQA and Vicuna’s TQA lies in the input embeddings at the 0th layer. In Vicuna, all input embeddings are word embeddings, whereas in Llava, the inputs are a mix of image embeddings and word embeddings (see Figure 1 a and c). In Vicuna, the color position contains color information, and the animal position contains animal information. When the color position’s word embedding is projected into the embedding matrix $E$, the color token ranks first. A crucial question arises: Can this method be applied to analyze information stored in visual embeddings? Specifically, will the color and animal tokens rank highly when the visual embeddings are projected into the embedding matrix $E$? To investigate this, we projected the top 20 positions’ visual embeddings into embedding space $E$ and computed the MRR scores for different tokens. The MRR for the correct color versus a random color is 0.455 versus 0.013, and for the correct animal versus a random animal, it is 0.076 versus 0.0003. We also calculated the MRR for the correct color and correct animal at a random position, which were 0.003 and 0.004, respectively. These results suggest that the top 20 positions already contain substantial information about the correct animal and color, whereas the random position does not.

### v2

@@ Abstract
Understanding the mechanisms behind Large Language Models (LLMs) is crucial for designing improved models and strategies. While recent studies have yielded valuable insights into the mechanisms of textual LLMs, the mechanisms of Multi-modal Large Language Models (MLLMs) remain underexplored. In this paper, we apply mechanistic interpretability methods to analyze the visual question answering (VQA) mechanisms in Llava. We compare the mechanisms between VQA and textual QA (TQA) in color answering tasks and find that: a) VQA exhibits a mechanism similar to the in-context learning mechanism observed in TQA; b) the visual features exhibit significant interpretability when projecting the visual embeddings into the embedding space; and c) Llava enhances the existing capabilities of the corresponding textual LLM Vicuna during visual instruction tuning. Based on these findings, we develop an interpretability tool to help users and researchers identify important visual locations for final predictions, aiding in the understanding of visual hallucination. Our method demonstrates faster and more effective results compared to existing interpretability approaches. Code: https://github.com/zepingyu0512/llava-mechanism
@@ 3.2 Mechanism Exploration of Textual QA
Evidence c). Following Dar et al. (2022), we project the color position’s layer input vector $h_{p}^{l-1}$ into the unembedding space and calculate the MRR score for the animal tokens [A] and [A1]. In S0, the MRR for [A] is 0.756, while for [A1], it is 0.001, as shown in Figure 3(b). The logit difference between [A] and [A1] at the color position is 0.32. In S1, the MRR for [A] is 0.002, the MRR for [A1] is 0.715, and the logit difference between [A1] and [A] is 1.70. These results demonstrate that the layer input vector at the color position, particularly in the most important attention heads, effectively encodes the animal features present in its context.
@@ 3.3 Mechanism Exploration of Visual QA
Evidence c) and d). We project the layer inputs of the top 20 important positions into unembedding space and compute the MRR and logit differences between the correct animal and a different animal. The correct animal’s MRR score is 0.318, while the other animal’s MRR score is 0.0004, as shown in Figure 4(b). The logit difference is 1.53, confirming that the important heads’ layer inputs contain crucial information about the animals. Furthermore, when the animal in the question is replaced with another animal, the attention score at the top 20 positions drops significantly from 0.807 to 0.564 (see Figure 4c), indicating that the last position encodes information about the question.
@@
Evidence e). A key difference between Llava’s VQA and Vicuna’s TQA lies in the input embeddings at the 0th layer. Vicuna uses only word embeddings, while Llava combines image and word embeddings (see Figure 1a and 1c). In Vicuna, the color and animal positions encode respective information, with the color token ranking first when projected into the embedding matrix E. To analyze visual embeddings, we projected the top 20 positions into E and computed MRR scores. For the correct color versus a random color, the MRR was 0.455 versus 0.013, and for the correct animal versus a random animal, 0.076 versus 0.0003. At random positions, the MRR scores for the correct color and animal were 0.003 and 0.004, respectively. These results suggest the top 20 positions encode significant information about the correct animal and color, while random positions do not.
@@
Results on other questions. To examine whether the mechanism applies to other questions, we replace the original question, ”What is the color of the [animal]?”, with ”What is the animal in this picture?” for comparison. Our findings show that the mechanisms are similar. Analysis of the identified important image patches reveals that these patches are closely associated with the animals. The sum of attention scores on the animal patches is 0.74, indicating that a substantial amount of information is extracted from these patches. Several examples are provided in Appendix A, and additional examples can be explored using the interpretability tool introduced in Section 4.